### v3

# Multi-Task Models Adversarial Attacks

Abstract
Multi-Task Learning (MTL) involves developing a singular model, known as a multi-task model, to concurrently perform multiple tasks. While the security of single-task models has been thoroughly studied, multi-task models pose several critical security questions, such as 1) their vulnerability to single-task adversarial attacks, 2) the possibility of designing attacks that target multiple tasks, and 3) the impact of task sharing and adversarial training on their resilience to such attacks. This paper addresses these queries through detailed analysis and rigorous experimentation. First, we explore the adaptation of single-task white-box attacks to multi-task models and identify their limitations. We then introduce a novel attack framework, the Gradient Balancing Multi-Task Attack (GB-MTA), which treats attacking a multi-task model as an optimization problem. This problem, based on averaged relative loss change across tasks, is approximated as an integer linear programming problem. Extensive evaluations on MTL benchmarks, NYUv2 and Tiny-Taxonomy, demonstrate GB-MTA’s effectiveness against both standard and adversarially trained multi-task models. The results also highlight a trade-off between task accuracy improvement via parameter sharing and increased model vulnerability due to enhanced attack transferability.
Introduction
Multi-task learning (MTL) leverages a single machine learning model to simultaneously address multiple tasks, such as semantic segmentation, depth estimation, and other vision-based prediction tasks. The single model, called multi-task model, shares parameters between tasks and is shown to have lower inference costs and higher generalization performance compared to single-task models without parameter sharing (Ruder 2017; Zhang, Liu, and Guan 2022a; Yao et al. 2020). Multi-task models are widely used in practical applications with high-security requirements, such as robotics and autonomous driving (Leang et al. 2020; Kokkinos 2017; Arcari et al. 2023). In this paper, we focus on the branched multi-task models, which are the most representative in the MTL literature (Zhang, Liu, and Guan 2022b). We use the terms task sharing and parameter sharing interchangeably.
In parallel to the developments in MTL, the security of single task classifiers has been put into question due to the existence of adversarial examples (Goodfellow, Shlens, and Szegedy 2014). An adversarial example is an input to a machine learning model (typically an image) that has been manipulated such that the model misclassifies the example with high confidence, but a human can still correctly recognize the input. Adversarial examples can be generated through white-box or black-box attacks depending on the assumed capabilities of the attacker (Tramer et al. 2020; Mahmood et al. 2021a). White-box attacks are generally considered more powerful (Carlini et al. 2019) because the attacker has access to the parameters and structure of the trained model.
Although adversarial attacks have been extensively studied on single-task models (Liu et al. 2016; Mahmood et al. 2022; Xu et al. 2022), related work on multi-task models is scarce. A pioneering study (Mao et al. 2020) pointed out that the adversarial robustness of deep neural networks increases as the number of tasks increases. MTA (Guo et al. 2020) tries to develop attacks in the MTL setting; however, the adversarial samples generated are task-specific and thus fail to attack all tasks simultaneously. Some other works (Gurulingan, Arani, and Zonooz 2021; Sobh et al. 2021) attempted to attack multi-task models by generating adversarial examples for each image while attacking one task at a time. Several critical security research questions (RQ) on MTL remain unclear:
RQ1: How secure are multi-task models to conventional single task adversarial attacks?
RQ2: Can adversarial attacks be designed to attack multiple tasks simultaneously?
RQ3: Does task sharing and adversarial training increase multi-task model robustness to adversarial attacks?
This paper answers the three questions through careful analysis and rigorous experimentation. To answer RQ1, we develop two naïve adaptations of single task white-box attacks for multi-task models, and analyze their inherent drawbacks. To answer RQ2, we propose a novel attack framework, GB-MTA (Gradient Balancing Multi-Task Attacker) to generate adversarial samples effective in attacking all tasks in a multi-task model. GB-MTA frames the problem of finding a unified attack perturbation in MTL as an optimization problem based on the averaged relative loss change (Sun et al. 2019; Zhang, Liu, and Guan 2022a) across tasks and solves the problem by approximating it as an Integer Linear Programming (ILP). To answer RQ3, we experiment with different levels of task sharing and demonstrate that there is a fundamental trade-off: Improving task accuracy and model efficiency through parameter sharing can increase the model’s vulnerability to adversarial attacks designated for these related tasks. We further explore the defense side of MTL, by adversarially training models with examples generated by GB-MTA. Our contributions to advancing the security of the field of MTL are summarized as follows.
Dynamic Gradient Balancing Multi-task Attack Framework - We formulate the MTL adversarial attack as an optimization problem with a specially designed multi-task objective function. To solve this optimization problem, we introduce a novel approach, GB-MTA that balances gradients from multiple tasks in a multi-task model when creating adversarial samples that work across all tasks.
Multi-Task Models Robustness Trade-off - We empirically demonstrate that task sharing can undermine model robustness due to increased attack transferability. In NYUv2, as the level of task sharing increases, the task attack transferability increases by , , and , respectively, when attacking segmentation, depth estimation, or normal prediction.
Multi-Task Learning Adversarial Training - We incorporate adversarial examples into the training of multi-task models to defend against adversarial attacks. The robustness of the models improves markedly, measured by the decreased accuracy drop after attacking from to . When attacking these adversarially trained models, GB-MTA still outperforms baselines by up to .
Attack Framework
This section first discusses existing white-box attacks and our adversarial threat model. This section then shows how single task white-box attacks can be adapted to multi-task models. We formulate the GB-MTA framework in the next section.
Single Task White-Box Attacks
In general, adversarial attacks can be formulated as follows (Madry et al. 2018). Let represent a clean input and its corresponding label. An attacker adds an adversarial perturbation to the input , to maximize the value of a loss function :
where denotes the parameters of the trained model under attack, and represents the maximum amount the adversary can perturb the input according to a given -norm. For notational simplicity, we omit in our future derivations.
Threat Model: In this paper, we focus on the untargeted white-box adversarial threat model (Carlini et al. 2019) as this represents one of the strongest and most widely used adversarial machine learning formulations (Carlini and Wagner 2017; Dong et al. 2018; Croce and Hein 2020b). In this setup, the attacker has knowledge of the model structure, trained model parameters and the corresponding loss function . In terms of bounds on the adversarial perturbation, we use one of the most widely used norms, , in line with previous works (Guo et al. 2020; Xu et al. 2022; Rathbun et al. 2022).
Single Task Attacks: In the white-box setting, one of the most prevalent strategies for generating the adversarial perturbation is to maximize the loss function by following the gradient ascent direction. This was originally done with the Fast Gradient Sign Method (FGSM) attack proposed in (Goodfellow, Shlens, and Szegedy 2014). Since the advent of FGSM, numerous improvements to the attack have been proposed. Although enumerating all the improvements in FGSM is beyond the scope of this paper, several important attack updates are worth noting. Updated attacks include the Projected Gradient Descent (PGD) attack (Madry et al. 2018), which adds a randomized start and makes FGSM iterative. The Momentum Iterative Momentum (MIM) (Dong et al. 2018) adds momentum to the gradient ascent optimization. More recently, in APGD (Croce and Hein 2020b), an adaptive step size has been shown to be one of the most effective white-box attacks for a single task, even against adversarial trained models (Mahmood, Mahmood, and Van Dijk 2021).
Naïve Multi-Task Attacks
RQ1: How secure are multi-task models to conventional single task adversarial attacks? This subsection answers the question by presenting two strategies for adapting single task white-box attacks to the multi-task formulation of the problem. It then discusses their inherent flaws, which are further demonstrated empirically. We denote these two attacks as naïve multi-task attacks.
In the case of a single task, untargeted white-box attack, an adversarial example can be generated (Madry et al. 2018) iteratively:
where represents the perturbation function associated with a specific white-box attack, represents the loss function for a single task, is the magnitude of the perturbation added in the current iteration of the attack and . Lastly, is the projection operation (Croce and Hein 2020b) to bound the adversarial sample within a specified range. In MTL, each input is associated with a set of true labels for tasks . Each task has its own task-specific loss function .
Single Attack - The first way in which multi-task models can be attacked is by focusing on only a single task’s gradient and ignoring the gradients of the rest tasks (Sobh et al. 2021; Gurulingan, Arani, and Zonooz 2021). For example, from Eq. 2, APGD (Croce and Hein 2020b) can be adapted to attack a single task :
where is a hyperparameter in APGD that controls the influence of previous update steps on the current update step. is the objective function of the task being attacked. In the experiment, we use Single-x to represent performing Single attack on a specific task x.
Total Attack - The second way in which single task attacks can be converted to attack multi-task models is through totaling all associated task loss functions , via summation. For example, Eq. 2 with single task Projected Gradient Descent (PGD) (Madry et al. 2018) can be adapted for Total-PGD:
We show the different attack formulations for Single-x with APGD and Total with PGD, but it is important to note that any combination of existing white-box attacks and adaptations can be made.
Naïve Attack Limitations: Both the Single and Total attacks come with significant drawbacks. The effectiveness of the Single attack is based on the following assumption: using one task’s attack direction will guarantee attack success on all other tasks. However, this assumption does not always hold, as the effectiveness of a Single attack is restricted due to the limited transferability of attack methods (Mahmood, Mahmood, and Van Dijk 2021). For example, in Figure 1, we show that when attacking the segmentation or the depth estimation task solely (Single-Segm or Single-Depth), the effectiveness of the attack on the normal prediction task (Norm ARP) is limited. The definition of ARP is elaborated in the Experiments Section.
Likewise, the effectiveness of the Total attack is based on an underlying assumption: for an adversarial example to work across all tasks, no one task’s gradient should dominate to avoid the limited attack transferability issue that Single faces. We denote it as the non-dominant magnitude assumption. However, the issue of gradient dominance is well recognized within the MTL literature and has garnered significant attention in the field of MTL optimizers (Yu et al. 2020; Navon et al. 2022). Empirically, we observe from Figure 1 that the Total attack exhibits a pattern similar to Single-Segm, indicating that the segmentation task dominates the gradient directions.
We also consider a modified version of Total where we take the sign of the gradients before the summation. In this way, the non-dominant magnitude assumption can be circumvented. We denote this attack as SignTotal. However, we empirically show that this attack is also not effective in Section Results on Attack Performance, as completely ignoring task’s gradient magnitudes also leads to a suboptimal attack.
The limitations due to the underlying assumptions of the Total and Single attacks mandate the need for an attack method tailored to multi-task models. The adversarial samples constructed from Single attacks are task-specific, and thus are not effective on non-targeted tasks. On the other hand, although the adversarial samples created from Total attack are task-agnostic, they are effective on only the tasks whose gradients dominate in MTL. An effective multi-task attack should be able to generate task-agnostic adversarial samples that are effective on all tasks in a multi-task model.
GB-MTA Framework
RQ2: Can adversarial attacks be designed to attack multiple tasks simultaneously? Dynamic Gradient Balancing Multi-task Attack (GB-MTA) builds on the success of existing single task adversarial attacks, while addressing the challenge in attacking multi-task models. GB-MTA accomplishes this by actively balancing the gradients across tasks, to derive an adversarial perturbation that is effective on all tasks. In this section, we first formulate a new attack optimization problem tailored for multi-task models. Since the problem is intractable, GB-MTA reformulates it to an Integer Linear Programming (ILP) problem, and then generates adversarial samples by solving the ILP problem.
Multi-Task Attack Optimization: We first reformulate the original single-task adversarial optimization introduced in Eq. 1 by decomposing . Here, represents the magnitude of the perturbation. represents the signed gradient direction vector, with values .
The above formulation is used to attack a single task. Attacking multiple tasks simultaneously in a multi-task model is fundamentally a multi-objective optimization problem. In this case, each objective function corresponds to one task. In adversarial machine learning, attacks are traditionally measured on a single task using one objective function that measures the attack success rate (Croce and Hein 2020b) or the robustness (Tramer et al. 2020). However, in MTL when evaluating two or more attacks, there is no single metric, as there are multiple tasks and each task has its own objective function value. This makes the comparison of two different attacks in MTL challenging.
Therefore, we formulate the multi-task attack optimization problem with a multi-task–specific objective function that aligns with the standard practice of assessing model performance in MTL. A multi-task model’s performance is typically measured by Average Relative Accuracy (ARA), denoted as , as opposed to using absolute values (Sun et al. 2019; Zhang, Liu, and Guan 2022a). The ARA metric compares the performance of a given multi-task model to that of a baseline model :
where is the number of tasks. is the average difference in accuracy between and in all tasks , normalized by the accuracy of . A higher indicates better model performance compared to the baseline.
When attacking a multi-task model, the goal is to substantially reduce the task performance of relative to , where now represents the model before the attack and denotes the model after the attack. will be a negative value, and the higher its absolute value is, the more effective the attack is. To find a perturbation direction that is the most effective, we reformulate the objective function in Eq. 5:
where represents the model loss difference for task before and after the attack to substitute the accuracy difference, since task loss in neural networks typically serves as a reliable indicator of task accuracy (i.e., higher task loss corresponds to lower task accuracy).
The optimization problem mentioned above can identify the optimal signed gradient direction vector but is intractable (Kreinovich, Lakeyev, and Noskov 1996; Horáček, Hladík, and Černỳ 2017). To address the problem, we apply the Taylor Expansion on , and reformulate it to an ILP problem as follows:
In practice, and are two matrices the same size as . Their product represents the dot product of the corresponding vectorized matrices.
Optimization Solution via Relaxed LP: GB-MTA identifies by first addressing a relaxed Linear Programming (LP) problem and then rounding the resulting solution to obtain an integer solution. In the first step, the LP relaxation will remove the requirement of integer values, i.e. , allowing them to be any real value instead. To solve the relaxed LP, we calculate the derivative of the objective function with respect to the variable as follows,
In other words, starts from the original state, that is, a zero matrix, and is updated along the direction of to maximize in the LP relaxation. Then in the second step, GB-MTA reintroduces the integer constraint, and the solution for the original ILP in Eq. 14 can be obtained by performing a rounding operation: .
The optimal suggests that the effective attack direction for multi-task models should be the sum of each task’s gradients dynamically weighted by its loss value. GB-MTA mitigates the dominating task issue in Total by dynamically balancing the gradients across tasks and avoids the limited transferability problem in Single-x by optimizing over all tasks simultaneously.
Integrating GB-MTA with Existing Attacks: Integrating GB-MTA with any existing attack can easily be accomplished, even for more advanced methods such as APGD (Croce and Hein 2020b). The key operation is to substitute the single task gradient, i.e., , with the balanced multi-task counterpart, i.e., . To illustrate. we provide the pseudocode for APGD integrated with GB-MTA in Algorithm C. The black text corresponds to the original APGD algorithm (Croce and Hein 2020b) and the part changed for GB-MTA is colored blue. Notice that on top of the key design shown in lines 2 and 11, we also change the absolute loss value used in the original APGD to the relative loss value sum over all the tasks in lines 4 and 14 to align with the MTL scenario.
Algorithm 1 GB-MTA-APGD
Experiments
Experimental Settings
Datasets and Tasks: We use two popular datasets in multi-task learning (MTL), NYUv2 (Silberman et al. 2012) and Tiny-Taskonomy (Zamir et al. 2018). The NYUv2 dataset consists of RGB-D indoor scenes and three tasks, 40-class semantic segmentation, depth estimation, and surface normal prediction. Tiny-Taskonomy contains RGB indoor images, and its five representative tasks are semantic segmentation, surface normal prediction, depth estimation, keypoint detection, and edge detection.
Evaluation Metrics and Loss Functions: Semantic segmentation uses a pixel-wise cross-entropy loss for each predicted class label. Surface normal prediction uses the inverse of cosine similarity between the normalized prediction and ground truth. All other tasks use the loss. Many tasks have distinct evaluation metrics with various scales. Hence, it is crucial to assess task performance in an equitable manner. Compounding the problem of different scales is the fact that some metrics are higher-the-better (e.g., accuracy, mean of intersection over union), while others are lower-the-better (e.g., distance, error). To address these issues and fairly measure the success of various attacks, we formulate a multi-task attack metric, Average Relative Performance (ARP):
where and represent the values of task ’s -th metric for the model before and after the attack respectively, and equals 0 if this metric is lower-the-better and 1 otherwise. denotes the number of metrics for task , and is the number of tasks. For each attack, we measure the corresponding ARP. A higher ARP indicates a higher performance drop and thus a more effective attack.
Multi-Task Models: We evaluate branched multi-task models from TreeMTL (Zhang, Liu, and Guan 2022b) using two backbone architectures: Deeplab-ResNet34 (Chen et al. 2017a) and MobileNetV2 (Sandler et al. 2018). We randomly sampled and trained 25 models with Deeplab-ResNet34 and 20 with MobileNetV2 for the NYUv2 dataset. For the Tiny-Taskonomy dataset, we sampled 15 models with Deeplab-ResNet34. These models cover a range of task sharing configurations from the all-shared models to those comprising an ensemble of independent single-task models.
Counterparts for Comparison: We compare GB-MTA to two types of baselines. The first type is the naïve multi-task attacks that repurpose existing single-task white-box attacks to multi-task models. It includes Total, SignTotal, and Single-x. We compare these baselines with GB-MTA by integrating them with three different single-task white-box attack methods, FGSM (Goodfellow, Shlens, and Szegedy 2014), PGD (Madry et al. 2018), and APGD (Croce and Hein 2020b). The second type of baseline is a multi-model attack method called Auto-SAGE, which is designed to attack multiple independent DNNs but can be directly applied to attack multi-task models. To compare fairly with Auto-SAGE, we integrate GB-MTA with Auto-SAGE instead of other single-task attacks.
Results on Attack Performance
Model Index | #Params (M) | PGD | AutoAttack | Auto-SAGE | |||||||||||
Segm | Norm | Dept | Total | SignTotal | GB-MTA | Segm | Norm | Dept | Total | SignTotal | GB-MTA | Baseline | GB-MTA | ||
IND | 63.83 | 28.18 | 18.23 | 62.14 | 50.52 | 67.85 | 73.41 | 30.74 | 23.52 | 75.62 | 58.30 | 69.29 | 87.58 | 71.07 | 88.53 |
5 | 62.48 | 29.12 | 30.00 | 70.99 | 61.40 | 71.46 | 79.60 | 32.06 | 39.25 | 88.03 | 70.34 | 73.05 | 94.44 | 84.51 | 95.70 |
35 | 62.25 | 36.46 | 32.54 | 98.11 | 68.30 | 91.24 | 100.53 | 41.01 | 41.71 | 119.52 | 78.10 | 95.22 | 121.30 | 124.97 | 131.56 |
41 | 61.13 | 38.06 | 46.02 | 100.41 | 67.51 | 95.20 | 103.03 | 42.09 | 59.54 | 120.39 | 75.42 | 99.05 | 122.50 | 113.15 | 126.94 |
21 | 55.65 | 31.45 | 39.55 | 79.10 | 65.00 | 76.70 | 84.81 | 34.80 | 57.48 | 96.28 | 74.13 | 78.75 | 100.94 | 92.91 | 102.72 |
39 | 55.43 | 36.03 | 45.68 | 83.78 | 67.83 | 90.40 | 96.63 | 39.81 | 56.61 | 106.07 | 76.16 | 94.08 | 115.62 | 108.61 | 118.71 |
26 | 42.54 | 34.19 | 45.39 | 80.94 | 69.39 | 88.58 | 93.88 | 38.32 | 58.72 | 101.69 | 80.18 | 92.38 | 115.79 | 107.18 | 118.67 |
AS | 21.28 | 46.65 | 53.36 | 105.74 | 58.79 | 83.56 | 88.51 | 56.39 | 69.60 | 133.54 | 67.42 | 88.10 | 108.57 | 112.56 | 119.22 |
Model Index | #Params (M) | PGD | AutoAttack | Auto-SAGE | |||||||||||||
Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | Baseline | GB-MTA | ||
IND | 106.38 | 248.04 | 23.13 | 117.44 | 12.48 | 11.50 | 282.01 | 274.17 | 245.25 | 23.21 | 117.83 | 12.40 | 11.37 | 279.28 | 270.25 | 282.01 | 274.17 |
190 | 105.03 | 190.91 | 49.41 | 138.67 | 14.98 | 14.02 | 232.55 | 236.63 | 175.02 | 45.06 | 130.69 | 14.26 | 13.64 | 213.47 | 218.29 | 216.32 | 218.88 |
358 | 103.68 | 189.90 | 50.42 | 152.70 | 18.14 | 15.53 | 237.12 | 247.29 | 174.80 | 46.32 | 142.73 | 17.07 | 14.73 | 217.99 | 227.61 | 220.93 | 228.97 |
959 | 96.86 | 201.59 | 48.17 | 163.56 | 15.56 | 17.64 | 241.75 | 256.35 | 183.75 | 44.46 | 150.37 | 14.70 | 17.04 | 221.31 | 234.58 | 223.54 | 234.56 |
1020 | 83.53 | 222.92 | 49.98 | 139.97 | 18.25 | 18.01 | 259.27 | 263.03 | 202.50 | 44.34 | 130.95 | 17.02 | 16.89 | 235.80 | 240.26 | 238.95 | 241.14 |
1043 | 75.59 | 205.27 | 44.26 | 145.47 | 23.99 | 38.96 | 246.90 | 250.79 | 189.32 | 39.97 | 136.29 | 22.55 | 36.50 | 227.23 | 233.25 | 230.54 | 234.28 |
1037 | 62.48 | 216.15 | 45.51 | 152.47 | 17.61 | 47.35 | 252.55 | 261.82 | 197.99 | 41.05 | 141.10 | 16.88 | 40.50 | 230.86 | 240.70 | 240.00 | 246.43 |
AS | 21.28 | 231.23 | 77.53 | 104.10 | 16.02 | 22.40 | 231.08 | 196.07 | 227.99 | 76.06 | 103.84 | 15.97 | 21.87 | 226.97 | 192.85 | 231.07 | 196.07 |
This subsection compares GB-MTA with baselines on their effectiveness in attacking multi-task models. Tables 1 and 2 compare the attack performance of the baselines and GB-MTA at the model level on NYUv2 and Tiny-Taskonomy respectively at . Overall, GB-MTA integrated with PGD and APGD have the highest ARP in 7 out of the 8 models on NYUv2, and in 6 out of the 8 models on Tiny-Taskonomy.
GB-MTA outperforms baselines Total, SignTotal, and Single-x in attack performance because it alleviates the baselines’ limitations discussed in Attack Framework. Single-x achieves limited attack effectiveness on tasks that are not the attack target x; Total shares a similar pattern of attack effectiveness across tasks as Single-Segm due to the issue of gradient dominance, making it less effective in attacking all tasks simultaneously. In contrast, GB-MTA dynamically balances the attack directions for all tasks, making it more threatening for systems that require high robustness. This is also the reason why GB-MTA outperforms the multi-model attack approach AutoSAGE — GB-MTA balances the gradients in AutoSAGE.
We further vary the maximum perturbation bound from 1 to 16 and report the average of the ARP over all multi-task models with diverse architectures in Figure 2 (Deeplab-ResNet34) and Figure 5 (MobileNet). We denote the average ARP as the overall ARP. GB-MTA is the best-performing attack in almost all and dataset settings, and when integrated with any of the three existing single-task (i.e., FGSM, PGD, APGD) or the multi-model attack (i.e., Auto-SAGE) approaches. There are a few cases, where for , GB-MTA and Total converge or perform almost identically. This is because at higher values, the magnitude of the noise becomes larger (and thus more visible) and all attacks become more effective.
Results on Attack Transferability
RQ3.1: Does task sharing increase the robustness of multi-task models to adversarial attacks? Existing literature in multi-task learning determines the appropriate level of parameter sharing with a focus on optimizing task accuracy. The results reported in this section, however, reveal a fundamental trade-off between the improvement in task accuracy due to positive task interactions and the increased vulnerability to adversarial attack due to the greater transferability of attacks from parameter sharing. We observe that a higher degree of parameter sharing between correlated tasks is associated with increased attack transferability.
We first define attack transferability in the MTL context. When using an attack method like Single-x, which attacks only one task in a multi-task model, the targeted task x, would be impacted the most. Referring to Figure 1, we can see that this occurs, since the highest bar for each of the Single-x attacks is the targeted task x. Therefore, we consider the degradation of the performance of the targeted task x, represented by ARP-x, as the upper bound for the attack effectiveness of Single-x. Since adversarial examples designed to attack one task may also be adversarial (e.g. misclassified) for another task, we can quantify this phenomenon in the MTL domain. Let x represent a task under attack, and y represent another task. The transferability of attack Single-x is
where represents the number of tasks. In short, we measure transferability as a performance degradation ratio. This ratio is the performance of the attack on all non-targeted tasks versus the performance of the attack on the targeted task.
Figure 4 illustrates the relationship between the levels of task sharing and attack transferability, showcasing the results for six multi-task models attacked by APGD Single-x variants. These models represent six levels of parameter sharing, ranging from all-share (AS/5L), where all layers of the backbone model (5L) are shared, to independent models (IND/0L), where no layers are shared. We observe a positive correlation between the degree of task sharing and the transferability of attacks in the three Single-x variants. As the level of task sharing increases, attack transferability also increases, from to () for Single-Dept, to () for Single-Segm, and to () for Single-Norm. These findings suggest a trade-off in multi-task model design: While sharing more parameters among related tasks can enhance task accuracy, it also amplifies attack transferability, thereby reducing model robustness, even when facing single task attacks. This underlines the importance of balancing accuracy and robustness in multi-task model design.
Attack on Adversarially Trained Multi-Task Models
RQ3.2: Does adversarial training increase MTL model robustness to adversarial attacks? A common single task strategy to defend against adversarial attacks is to leverage adversarial training (Madry et al. 2018; Bai et al. 2021; Zhang et al. 2020). This defense involves generating adversarial samples and using them as part of the dataset during training. This typically results in reduced model performance on clean inputs, but increased robustness to adversarial attacks. To the best of our knowledge, no existing work has dealt with the open question of whether adversarial training can be applied effectively to multi-task models.
Adv. Train | Single-Segm | Single-Norm | Single-Dept | Total | GB-MTA |
w/o AT | 46.65 | 53.36 | 105.74 | 58.79 | 88.51 |
Single-Segm | 14.67 | 6.60 | 16.70 | 19.41 | 19.92 |
Single-Norm | 20.20 | 10.61 | 25.52 | 25.78 | 29.26 |
Single-Dept | 16.60 | 8.41 | 14.12 | 18.48 | 20.31 |
Total | 13.21 | 6.26 | 13.16 | 16.03 | 16.62 |
GB-MTA | 13.58 | 5.97 | 12.76 | 15.98 | 16.64 |
We adopt the single task Friendly Adversarial Training (FAT) (Zhang et al. 2020) to MTL. FAT is a PGD-based adversarial training method. In order to apply this defense effectively to MTL, we modify the default PGD to GB-MTA-PGD to use adversarial examples generated from multiple tasks as opposed to just single task adversarial examples. Our new defense is coined FAT GB-MTA. We also train multi-task models using FAT with the naïve MTL attacks Single and Total. After adversarial training, we re-evaluate the multi-task attacks to assess both model robustness and attack effectiveness.
Table 3 reports the ARP of 30 cases, i.e., (without adversarial training + 5 adversarially trained models) 5 attack methods. We make two main observations. First, the notable reduction in ARP of attack methods (e.g., from to when attacking with Single-Dept) demonstrates that models enhanced with adversarial training are substantially more robust than the model without such training. Second, from the perspective of attack performance, GB-MTA is still the most effective attack method, consistently outperforming the Single-x and Total baselines by up to .
Conclusion
In this work, we make significant developments in multi-task learning (MTL) security through novel attack design, empirical exploration with multi-task models and robustness analyses. We first analyzed naïve adaptions of single task white-box attacks to the MTL domain and experimentally demonstrated their ineffectiveness. We then developed a new framework, Dynamic Gradient Balancing Multi-task Attack (GB-MTA), to effectively attack all tasks in multi-task models. On models trained on the NYUv2 and Tiny-Taskonomy datasets, GB-MTA achieves the highest overall attack strength and is the strongest attack on 6 out of 8 models for both datasets. We further analyzed the adversarial transferability of MTL adversarial examples and discovered a new phenomenon: Task sharing can lead to increased adversarial transferability. Lastly, from the defense side, we adversarially trained multi-task models in a new approach, which we coined FAT GB-MTA. GB-MTA is the most effective attack, even on these multi-task models, showcasing its effectiveness as a benchmark for analyzing future MTL defenses.
Acknowledgement
This material is based upon work supported by the National Science Foundation under Grant No. 2312396, 2220211, 2224054, and 2247893. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
Appendix A Related Works
Adversarial Attacks. Adversarial attacks fall into two main categories: white-box and black-box attacks (Mahmood et al. 2021b). In white-box attacks, an attacker has access to the target model’s internal information, enabling direct gradient extraction and adversarial example generation (Carlini et al. 2019). On the contrary, in black-box attacks, an attacker has limited model knowledge and uses alternative information sources (Chen et al. 2017b; Zhou et al. 2020) to create adversarial examples. This paper focuses on white-box attacks, as they are generally more effective compared to black-box attacks (Croce and Hein 2020b; Wang et al. 2022).
In recent years, many white-box attack techniques have been developed. Fast Gradient Sign Method (FGSM) (Goodfellow, Shlens, and Szegedy 2014) generates adversarial examples by introducing non-random noise in the gradient direction of the loss function. Projected Gradient Descent (PGD) (Madry et al. 2018) and Momentum Iterative Method (MIM) (Dong et al. 2018) improve FGSM by generating adversarial samples in an iterative process. Later, Croce and Hein (Croce and Hein 2020b) proposed Auto Projected Gradient Descent (APGD) with adaptive step size and combined it with two complementary attacks (Croce and Hein 2020a; Andriushchenko et al. 2020), developing the ensemble method APGD that outperforms existing methods on diverse benchmark datasets. In addition to gradient-based strategies, alternative methods have emerged. For instance, the Backward Pass Differentiable Approximation (BPDA) (Athalye, Carlini, and Wagner 2018) accommodates non-differentiable functions, while the Carlini and Wagner () attack (Carlini and Wagner 2017) perturbs images with minimal delta to misclassify them.
Multi-Task Learning. In Multi-Task Learning (MTL), researchers develop memory and computation-efficient multi-task models that simultaneously address multiple tasks (Ruder 2017; Yao et al. 2020). The main challenge lies in determining the parameters to share across tasks to optimize both resource efficiency and task accuracy. This has led to an abundance of multi-task model architectures designed either manually (Huang et al. 2015; Jou and Chang 2016; Dvornik et al. 2017; Kokkinos 2017; Ranjan, Patel, and Chellappa 2017) or automatically (Sun et al. 2019; Zhang, Liu, and Guan 2022b; Ahn, Kim, and Oh 2019; Guo, Lee, and Ulbricht 2020; Zhang, Liu, and Guan 2022a). This paper focuses on attacking branched multi-task models, which are the most representative in the MTL literature. Besides, branched multi-task models have a wide range of sharing patterns across tasks, facilitating the study of the relationship between the robustness of multi-task models and their sharing patterns.
Robustness of Multi-Task Models. Few studies have examined the robustness of the model in MTL settings. A pioneering study (Mao et al. 2020) pointed out that the adversarial robustness of deep neural networks increases as the number of tasks increases. Subsequent research (Klingner, Bar, and Fingscheidt 2020; Ghamizi et al. 2022) further emphasizes the importance of selecting suitable tasks for joint learning to create more robust models. While these studies offer intriguing insights, they do not specifically propose adversarial attack methods for multi-task models. Regarding attacks on multi-task models, MTA (Guo et al. 2020) tries to develop attacks in the MTL setting, however, the generated adversarial samples are task-specific and thus fail to attack all the tasks simultaneously. Some other work (Gurulingan, Arani, and Zonooz 2021; Sobh et al. 2021) attempted to attack the multi-task model by generating adversarial examples for each image while attacking one task at a time.
Appendix B MTL Attack Optimization Approximation
The optimization problem we formulate for multi-task attack in Section 3 of the main paper is,
Here, represents the model loss difference for task before and after the attack.
As the problem is intractable, we make some approximations and reformulate it be an Integer Linear Programming (ILP) problem. To do so, we first apply the Taylor expansion on in the numerator of the objective function at the point of :
We ignore the remainder because, in the context of adversary attacks, we have , indicating that the change is sufficiently small.
The proposed approximate optimization problem for multi-task attacks is thus formulated as follows:
where and are constants that can be ignored when solving the optimization problem. In practice, and are two matrices with the same size as , thus their product represents the dot product of the corresponding vectorized matrices.
Appendix C Algorithm Pseudocode
As presented in Section 3 in the main paper, integrating GB-MTA with any existing attack can be easily accomplished by substituting the single-task gradient, i.e., , with the balanced multi-task counterpart, i.e., . We illustrate how to integrate GB-MTA into a single-task attack algorithm APGD and a multi-model attack algorithm Auto-SAGE.
APGD and GB-MTA-APGD
We provide pseudocode comparisons for APGD with and without integrating GB-MTA in Algorithms C and C. We color the difference in the two algorithms in blue. To integrate GB-MTA in APGD, we first change the key part of the code determining the attack direction (lines 2 and 11) to the proposed balanced multi-task gradients. Then we further update the objective function (lines 4 and 14) from the absolute loss value to the relative loss value sum over all tasks to accommodate the multi-task setting. All other lines in Algorithm C are kept the same as the original APGD.
Algorithm 2 APGD
Algorithm 3 GB-MTA-APGD
Auto-SAGE and GB-MTA-Auto-SAGE
For Auto-SAGE (Rathbun et al. 2022), the original attack formulation is
To integrate GB-MTA with Auto-SAGE, we normalize the gradient of the objective function and with the objective function value. The updated attack will be,
Appendix D More Experimental Results
This section reports evaluation metrics and more experimental results that are ommited from the main paper.
Evaluation Metrics
Semantic segmentation is evaluated using mean Intersection over Union and Pixel Accuracy (mIoU and Pixel Acc, the higher the better) in NYUv2.
Surface normal prediction is evaluated using mean and median angle distances between the prediction and the ground truth (the lower the better), and the percentage of pixels whose prediction is within the angles of , and to the ground truth (the higher the better).
Depth estimation uses the absolute and relative errors between prediction and ground truth (the lower the better). Furthermore, the percentage of pixels whose prediction is within the thresholds of to the ground truth, i.e. , is used (the higher the better).
Tiny-Taskonomy is evaluated using the task-specific loss of each task directly.
Results of Attack Performance
Figure 5 illustrates the attack performance on NYUv2 with MobileNetV2. To be consistent with the main paper, we conduct the experiments using different variants of GB-MTA and the naïve multi-task attacks. The x-axis represents the perturbation bound ranging from 1 to 16, while the y-axis displays the overall Average Relative Performance (ARP, higher-the-better). Overall, GB-MTA consistently outperforms baselines.
We present the full tables of ARP after the attack of all 25 multi-task models trained on NYUv2 with Deeplab-ResNet34 for perturbation bound in Tables 4 and 5. Similarly, Tables 6 and 7 show the attack results for all 15 multi-task models for Tiny-Taskonomy. Overall, GB-MTA achieves 80% first place (80 out of 100 cases) on NYUv2 and 88.33% (53 out of 60) on Tiny-Taskonomy, demonstrating the effectiveness of adversarial samples from GB-MTA.
We also show the evaluation results with perturbation bound in Tables 8 and 9 for NYUv2 and Tables 10 and 11 for Taskonomy.
Model Index | #Params (M) | FGSM | PGD | ||||||||||
Segm | Norm | Dept | Total | SignTotal | GB-MTA | Segm | Norm | Dept | Total | SignTotal | GB-MTA | ||
IND | 63.83 | 21.56 | 16.97 | 35.24 | 33.47 | 38.1 | 39.07 | 28.18 | 18.23 | 62.14 | 50.52 | 67.85 | 73.41 |
14 | 63.6 | 21.72 | 27.07 | 35.96 | 38.51 | 46.54 | 46.69 | 29.54 | 44.67 | 63.07 | 63.09 | 92.95 | 98.9 |
4 | 63.6 | 21.12 | 17.09 | 39.01 | 38.32 | 43.45 | 44.22 | 27.92 | 21.74 | 63.88 | 59.03 | 72.88 | 79.05 |
9 | 63.6 | 26.2 | 21.63 | 38.79 | 37.72 | 40.91 | 43.21 | 33.56 | 19.35 | 83.54 | 70.83 | 85.73 | 92.65 |
23 | 63.59 | 26.18 | 23.64 | 40.69 | 37.52 | 41.52 | 43.8 | 33.2 | 26.47 | 80.66 | 65.38 | 84.67 | 90.43 |
5 | 62.48 | 23.19 | 22.05 | 39.14 | 38.54 | 39.67 | 42.99 | 29.12 | 30 | 70.99 | 61.4 | 71.46 | 79.6 |
10 | 62.48 | 26.1 | 19.72 | 39.59 | 34.76 | 39.99 | 41.57 | 37.65 | 21.7 | 97.24 | 60.91 | 86.69 | 93.55 |
28 | 62.47 | 26.12 | 28.27 | 40.01 | 38.67 | 42.65 | 44.93 | 33.36 | 49.64 | 89.3 | 72.06 | 101.31 | 107.59 |
35 | 62.25 | 25.12 | 24.02 | 37.27 | 33.53 | 36.93 | 39.73 | 36.46 | 32.54 | 98.11 | 68.3 | 91.24 | 100.53 |
38 | 62.25 | 25.94 | 26.21 | 35.87 | 34.49 | 37.22 | 39.52 | 34.42 | 43.02 | 83.04 | 63.46 | 87.27 | 93.76 |
41 | 61.13 | 26.08 | 26.38 | 40.88 | 36.12 | 39.79 | 42.64 | 38.06 | 46.02 | 100.41 | 67.51 | 95.2 | 103.03 |
11 | 55.66 | 28.75 | 20.65 | 39.53 | 35.69 | 39.38 | 41.37 | 38.52 | 18.99 | 93.81 | 57.74 | 82.58 | 89.68 |
21 | 55.65 | 25.27 | 23.64 | 42.66 | 39.93 | 40.26 | 44.65 | 31.45 | 39.55 | 79.1 | 65 | 76.7 | 84.81 |
33 | 55.43 | 27.67 | 25.99 | 38.89 | 36.26 | 36.67 | 40.35 | 33.67 | 35.08 | 87.47 | 65.02 | 77.33 | 87.08 |
36 | 55.43 | 27.8 | 29.84 | 37.76 | 34.08 | 39.49 | 41.36 | 38.36 | 40.88 | 96.69 | 57.78 | 91.97 | 99.72 |
39 | 55.43 | 29.97 | 30.95 | 40.68 | 39.95 | 41.62 | 44.76 | 36.03 | 45.68 | 83.78 | 67.83 | 90.4 | 96.63 |
44 | 54.32 | 25.99 | 25.9 | 39.41 | 34.13 | 37.99 | 40.86 | 40.48 | 43.64 | 104.05 | 63.19 | 93.5 | 100.56 |
42 | 54.32 | 28.27 | 29.47 | 38.55 | 36.15 | 39.21 | 41.42 | 36.9 | 52.1 | 91.33 | 60.92 | 89.17 | 95.79 |
48 | 47.5 | 29.01 | 28.54 | 38.05 | 34.44 | 36.5 | 38.38 | 39.88 | 48.02 | 92.03 | 58.34 | 84.81 | 90.34 |
26 | 42.54 | 25.47 | 23.76 | 38.8 | 37.75 | 38.87 | 41.97 | 34.19 | 45.39 | 80.94 | 69.39 | 88.58 | 93.88 |
17 | 42.54 | 25.5 | 25.78 | 41.79 | 39.93 | 39.61 | 44.12 | 32.21 | 41.36 | 79.58 | 66.8 | 74.34 | 84.75 |
34 | 42.32 | 27.11 | 24.73 | 35.06 | 31.75 | 34.69 | 36.48 | 41.27 | 30.49 | 86.88 | 52.77 | 77 | 82.03 |
43 | 41.21 | 28.1 | 27.35 | 39.21 | 34.18 | 37.55 | 39.68 | 42.74 | 44.65 | 104.83 | 58.03 | 90.72 | 95.97 |
49 | 34.39 | 28.23 | 28.02 | 39.85 | 34.31 | 37.66 | 39.55 | 44.96 | 60.65 | 119.52 | 59.28 | 98.39 | 106.54 |
AS | 21.28 | 29.5 | 28.31 | 38.91 | 34.71 | 36.66 | 39.4 | 46.65 | 53.36 | 105.74 | 58.79 | 83.56 | 88.51 |
Model Index | #Params (M) | AutoAttack | Auto-SAGE | ||||||
Segm | Norm | Dept | Total | SignTotal | GB-MTA | Baseline | GB-MTA | ||
IND | 63.83 | 30.74 | 23.52 | 75.62 | 58.3 | 69.29 | 87.58 | 71.07 | 88.53 |
14 | 63.6 | 32.43 | 52.73 | 80.01 | 73.77 | 95.69 | 118.63 | 94.14 | 118.97 |
4 | 63.6 | 30.62 | 30.53 | 76.81 | 67.7 | 73.99 | 93.73 | 77.63 | 94.41 |
9 | 63.6 | 37.53 | 25.86 | 100.49 | 80.34 | 88.5 | 108.6 | 102.27 | 110.24 |
23 | 63.59 | 37 | 32.9 | 98.26 | 73.23 | 87.13 | 107.08 | 102.43 | 110.87 |
5 | 62.48 | 32.06 | 39.25 | 88.03 | 70.34 | 73.05 | 94.44 | 84.51 | 95.7 |
10 | 62.48 | 41.62 | 28.69 | 122.36 | 67.27 | 89.72 | 113.78 | 110.51 | 119.37 |
28 | 62.47 | 37.28 | 59.37 | 113.04 | 83.65 | 105.58 | 130.93 | 118.33 | 133.16 |
35 | 62.25 | 41.01 | 41.71 | 119.52 | 78.1 | 95.22 | 121.3 | 124.97 | 131.56 |
38 | 62.25 | 38.3 | 54.74 | 106.5 | 71.85 | 90.27 | 113.04 | 105.82 | 117.23 |
41 | 61.13 | 42.09 | 59.54 | 120.39 | 75.42 | 99.05 | 122.5 | 113.15 | 126.94 |
11 | 55.66 | 43.1 | 24.76 | 112.77 | 63.12 | 84.97 | 106.21 | 107.85 | 115.3 |
21 | 55.65 | 34.8 | 57.48 | 96.28 | 74.13 | 78.75 | 100.94 | 92.91 | 102.72 |
33 | 55.43 | 37.37 | 44.63 | 110.04 | 74.07 | 79.81 | 104.87 | 109.74 | 112.81 |
36 | 55.43 | 43 | 50.84 | 117.13 | 63.6 | 95.66 | 119.98 | 112.12 | 127.46 |
39 | 55.43 | 39.81 | 56.61 | 106.07 | 76.16 | 94.08 | 115.62 | 108.61 | 118.71 |
44 | 54.32 | 45.29 | 54.94 | 128.01 | 68.93 | 92.54 | 120.52 | 125.6 | 130.93 |
42 | 54.32 | 41.34 | 67.03 | 116.01 | 67.46 | 92.41 | 114.74 | 113.28 | 120.97 |
48 | 47.5 | 44.42 | 61.56 | 111.9 | 62.92 | 84.19 | 106.66 | 107.63 | 114.49 |
26 | 42.54 | 38.32 | 58.72 | 101.69 | 80.18 | 92.38 | 115.79 | 107.18 | 118.67 |
17 | 42.54 | 35.91 | 51.57 | 98.39 | 75.5 | 76.19 | 98.4 | 100.51 | 103.25 |
34 | 42.32 | 45.56 | 38.38 | 108.03 | 55.82 | 79.52 | 96.55 | 95.69 | 105.38 |
43 | 41.21 | 47.44 | 59.7 | 131.97 | 61.11 | 89.67 | 115.1 | 122.87 | 128.99 |
49 | 34.39 | 48.4 | 70.91 | 136.53 | 61.73 | 92.81 | 119.81 | 125.91 | 130.8 |
AS | 21.28 | 56.39 | 69.6 | 133.54 | 67.42 | 88.1 | 108.57 | 112.56 | 119.22 |
Model Index | #Params (M) | FGSM | PGD | ||||||||||||
Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | ||
IND | 106.38 | 85.42 | 18.71 | 51.99 | 11.09 | 11.96 | 88.26 | 95.10 | 248.04 | 23.13 | 117.44 | 12.48 | 11.50 | 282.01 | 274.17 |
190 | 105.03 | 73.19 | 23.13 | 42.67 | 13.75 | 16.55 | 72.87 | 74.15 | 190.91 | 49.41 | 138.67 | 14.98 | 14.02 | 232.55 | 236.63 |
348 | 105.02 | 67.60 | 23.05 | 49.61 | 14.74 | 14.86 | 70.57 | 78.14 | 202.55 | 51.07 | 147.26 | 19.62 | 14.74 | 246.88 | 249.54 |
200 | 104.81 | 67.83 | 19.17 | 44.49 | 12.39 | 14.55 | 71.01 | 76.78 | 182.76 | 50.02 | 139.60 | 13.13 | 16.56 | 226.76 | 233.80 |
352 | 104.80 | 67.17 | 22.59 | 44.80 | 15.19 | 14.85 | 68.88 | 74.70 | 178.64 | 48.81 | 143.03 | 17.71 | 13.70 | 224.02 | 234.34 |
469 | 104.57 | 72.08 | 22.52 | 49.17 | 17.73 | 15.17 | 73.25 | 79.48 | 182.49 | 49.82 | 155.29 | 27.55 | 18.20 | 228.44 | 240.80 |
358 | 103.68 | 69.37 | 23.91 | 46.00 | 15.62 | 15.60 | 69.91 | 74.85 | 189.90 | 50.42 | 152.70 | 18.14 | 15.53 | 237.12 | 247.29 |
481 | 103.46 | 72.21 | 22.81 | 46.38 | 21.68 | 16.24 | 74.59 | 81.94 | 212.18 | 51.14 | 157.02 | 38.94 | 24.78 | 256.71 | 266.46 |
191 | 98.21 | 68.01 | 20.88 | 40.81 | 12.19 | 13.79 | 68.19 | 69.76 | 190.28 | 52.63 | 153.36 | 13.94 | 12.94 | 226.54 | 231.65 |
959 | 96.86 | 69.48 | 23.28 | 44.76 | 21.98 | 18.80 | 73.29 | 76.47 | 201.59 | 48.17 | 163.56 | 15.56 | 17.64 | 241.75 | 256.35 |
958 | 83.75 | 74.68 | 23.46 | 47.23 | 22.66 | 16.49 | 80.55 | 84.18 | 210.33 | 50.29 | 166.25 | 15.22 | 13.20 | 250.88 | 266.38 |
1020 | 83.53 | 74.85 | 22.43 | 48.32 | 20.34 | 16.31 | 80.78 | 86.30 | 222.92 | 49.98 | 139.97 | 18.25 | 18.01 | 259.27 | 263.03 |
1043 | 75.59 | 77.23 | 21.55 | 53.74 | 22.06 | 23.22 | 84.89 | 94.04 | 205.27 | 44.26 | 145.47 | 23.99 | 38.96 | 246.90 | 250.79 |
1037 | 62.48 | 72.67 | 22.63 | 48.99 | 22.62 | 17.82 | 77.62 | 84.68 | 216.15 | 45.51 | 152.47 | 17.61 | 47.35 | 252.55 | 261.82 |
AS | 21.28 | 76.86 | 45.37 | 54.92 | 21.83 | 21.60 | 76.24 | 76.03 | 231.23 | 77.53 | 104.10 | 16.02 | 22.40 | 231.08 | 196.07 |
Model Index | #Params (M) | AutoAttack | Auto-SAGE | |||||||
Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | Baseline | GB-MTA | ||
IND | 106.38 | 245.25 | 23.21 | 117.83 | 12.40 | 11.37 | 279.28 | 270.25 | 282.01 | 274.17 |
190 | 105.03 | 175.02 | 45.06 | 130.69 | 14.26 | 13.64 | 213.47 | 218.29 | 216.32 | 218.88 |
348 | 105.02 | 184.25 | 46.19 | 137.71 | 18.80 | 13.90 | 225.45 | 229.03 | 228.10 | 229.64 |
200 | 104.81 | 168.93 | 44.77 | 130.06 | 12.55 | 15.74 | 208.48 | 215.48 | 210.54 | 216.37 |
352 | 104.80 | 164.72 | 44.09 | 134.15 | 16.84 | 13.15 | 205.59 | 216.01 | 207.59 | 216.50 |
469 | 104.57 | 168.58 | 44.55 | 143.19 | 26.30 | 17.31 | 210.28 | 220.91 | 212.21 | 221.93 |
358 | 103.68 | 174.80 | 46.32 | 142.73 | 17.07 | 14.73 | 217.99 | 227.61 | 220.93 | 228.97 |
481 | 103.46 | 193.75 | 45.87 | 146.31 | 37.71 | 21.32 | 234.60 | 244.78 | 236.99 | 246.22 |
191 | 98.21 | 174.37 | 47.96 | 142.62 | 13.05 | 12.25 | 208.24 | 213.53 | 211.71 | 214.30 |
959 | 96.86 | 183.75 | 44.46 | 150.37 | 14.70 | 17.04 | 221.31 | 234.58 | 223.54 | 234.56 |
958 | 83.75 | 192.26 | 45.68 | 152.92 | 14.37 | 12.80 | 229.57 | 242.97 | 232.22 | 243.31 |
1020 | 83.53 | 202.50 | 44.34 | 130.95 | 17.02 | 16.89 | 235.80 | 240.26 | 238.95 | 241.14 |
1043 | 75.59 | 189.32 | 39.97 | 136.29 | 22.55 | 36.50 | 227.23 | 233.25 | 230.54 | 234.28 |
1037 | 62.48 | 197.99 | 41.05 | 141.10 | 16.88 | 40.50 | 230.86 | 240.70 | 240.00 | 246.43 |
AS | 21.28 | 227.99 | 76.06 | 103.84 | 15.97 | 21.87 | 226.97 | 192.85 | 231.07 | 196.07 |
Model Index | #Params (M) | FGSM | PGD | ||||||||||
Segm | Norm | Dept | Total | SignTotal | GB-MTA | Segm | Norm | Dept | Total | SignTotal | GB-MTA | ||
IND | 63.83 | 17.09 | 9.31 | 25.73 | 27.45 | 31.47 | 33.15 | 24.07 | 12.34 | 39.67 | 37.64 | 46.06 | 52.51 |
14 | 63.60 | 19.73 | 10.82 | 33.63 | 33.88 | 36.30 | 39.51 | 27.67 | 12.33 | 64.64 | 57.08 | 63.67 | 73.42 |
4 | 63.60 | 17.57 | 20.65 | 27.17 | 33.05 | 40.21 | 41.11 | 25.28 | 35.17 | 42.23 | 47.76 | 66.83 | 73.81 |
9 | 63.60 | 16.99 | 9.37 | 29.26 | 32.13 | 35.28 | 37.64 | 23.91 | 13.30 | 44.03 | 44.93 | 51.26 | 58.82 |
23 | 63.59 | 19.59 | 13.92 | 32.88 | 32.01 | 35.67 | 38.01 | 27.56 | 17.54 | 61.19 | 52.80 | 62.37 | 70.10 |
5 | 62.48 | 17.74 | 12.82 | 30.67 | 32.45 | 33.11 | 37.15 | 24.51 | 17.33 | 48.98 | 46.82 | 51.18 | 59.36 |
10 | 62.48 | 21.43 | 10.35 | 34.04 | 30.76 | 35.13 | 37.51 | 31.32 | 13.71 | 73.02 | 50.90 | 64.28 | 73.08 |
28 | 62.47 | 20.33 | 20.16 | 33.82 | 35.17 | 39.46 | 41.95 | 27.68 | 35.74 | 63.97 | 56.20 | 73.24 | 80.54 |
35 | 62.25 | 20.25 | 18.54 | 30.89 | 30.84 | 34.10 | 37.00 | 27.91 | 29.93 | 58.75 | 50.02 | 64.14 | 71.00 |
38 | 62.25 | 19.56 | 15.72 | 33.75 | 31.01 | 34.50 | 37.92 | 29.23 | 22.82 | 74.39 | 54.53 | 66.51 | 76.85 |
41 | 61.13 | 21.00 | 19.33 | 34.84 | 32.15 | 36.69 | 39.67 | 31.10 | 32.23 | 75.81 | 54.53 | 71.75 | 79.93 |
11 | 55.66 | 22.12 | 10.87 | 33.89 | 30.51 | 33.73 | 36.88 | 31.73 | 12.76 | 70.28 | 47.48 | 60.24 | 68.96 |
21 | 55.65 | 18.96 | 15.58 | 34.51 | 34.17 | 34.54 | 39.08 | 26.53 | 23.81 | 56.20 | 50.44 | 56.26 | 64.85 |
33 | 55.43 | 21.55 | 18.91 | 31.83 | 29.54 | 34.62 | 37.13 | 31.01 | 28.48 | 72.41 | 47.68 | 66.77 | 74.71 |
36 | 55.43 | 21.64 | 21.18 | 33.84 | 34.08 | 36.15 | 39.67 | 29.49 | 32.03 | 62.09 | 54.67 | 67.51 | 74.64 |
39 | 55.43 | 19.99 | 16.21 | 31.60 | 30.32 | 31.06 | 35.11 | 27.57 | 24.14 | 62.35 | 50.84 | 57.22 | 66.48 |
44 | 54.32 | 20.61 | 19.00 | 34.79 | 30.57 | 35.03 | 38.05 | 33.02 | 30.36 | 76.81 | 52.04 | 69.01 | 76.70 |
42 | 54.32 | 20.61 | 20.92 | 32.00 | 30.52 | 34.49 | 37.13 | 30.17 | 36.93 | 66.13 | 49.23 | 66.79 | 73.41 |
48 | 47.50 | 22.34 | 20.88 | 33.26 | 29.65 | 33.05 | 35.71 | 32.75 | 33.81 | 70.45 | 48.91 | 65.49 | 71.34 |
26 | 42.54 | 19.16 | 17.23 | 32.22 | 33.51 | 35.20 | 38.65 | 28.21 | 32.11 | 58.06 | 54.17 | 63.98 | 71.52 |
17 | 42.54 | 19.70 | 17.81 | 34.21 | 35.10 | 34.35 | 39.69 | 26.89 | 28.06 | 57.05 | 52.25 | 56.21 | 65.35 |
34 | 42.32 | 22.45 | 15.94 | 30.89 | 28.23 | 31.25 | 33.59 | 33.91 | 21.11 | 64.29 | 44.38 | 57.97 | 63.58 |
43 | 41.21 | 22.92 | 19.65 | 34.96 | 30.41 | 34.54 | 37.20 | 35.07 | 30.38 | 78.53 | 49.62 | 68.22 | 74.39 |
49 | 34.39 | 22.75 | 21.15 | 35.46 | 30.18 | 34.34 | 36.90 | 37.63 | 43.02 | 86.58 | 51.52 | 73.98 | 81.24 |
AS | 21.28 | 24.21 | 20.95 | 35.75 | 30.65 | 33.22 | 35.97 | 37.80 | 37.87 | 78.58 | 49.49 | 64.36 | 69.60 |
Model Index | #Params (M) | APGD | Auto-SAGE | ||||||
Segm | Norm | Dept | Total | SignTotal | GB-MTA | Baseline | GB-MTA | ||
IND | 63.83 | 27.16 | 15.28 | 48.07 | 43.27 | 46.68 | 62.72 | 50.60 | 63.78 |
14 | 63.60 | 31.53 | 15.75 | 77.52 | 65.34 | 64.94 | 86.42 | 80.49 | 88.43 |
4 | 63.60 | 28.65 | 41.38 | 51.30 | 54.93 | 67.88 | 88.39 | 66.46 | 87.21 |
9 | 63.60 | 27.10 | 17.16 | 52.03 | 51.10 | 51.84 | 69.28 | 58.22 | 70.20 |
23 | 63.59 | 31.31 | 21.70 | 74.76 | 59.86 | 63.49 | 83.53 | 78.71 | 85.70 |
5 | 62.48 | 27.91 | 21.99 | 59.31 | 54.01 | 51.87 | 70.46 | 62.59 | 71.73 |
10 | 62.48 | 35.73 | 17.45 | 91.60 | 57.20 | 65.46 | 88.04 | 81.73 | 92.48 |
28 | 62.47 | 31.55 | 43.78 | 80.85 | 65.45 | 74.95 | 98.42 | 84.99 | 97.51 |
35 | 62.25 | 32.18 | 37.56 | 72.78 | 57.23 | 65.31 | 85.37 | 74.98 | 85.89 |
38 | 62.25 | 34.16 | 28.68 | 92.92 | 63.05 | 68.05 | 94.19 | 92.69 | 99.99 |
41 | 61.13 | 35.40 | 40.66 | 93.92 | 62.34 | 73.43 | 96.80 | 86.64 | 100.73 |
11 | 55.66 | 36.47 | 16.01 | 87.30 | 53.81 | 61.46 | 83.26 | 80.89 | 88.73 |
21 | 55.65 | 29.81 | 32.22 | 68.68 | 57.18 | 57.15 | 76.41 | 69.18 | 78.34 |
33 | 55.43 | 35.83 | 35.63 | 90.57 | 54.05 | 68.32 | 92.08 | 83.96 | 96.00 |
36 | 55.43 | 33.51 | 39.54 | 77.22 | 61.78 | 68.97 | 88.82 | 80.06 | 89.48 |
39 | 55.43 | 31.24 | 29.83 | 78.41 | 58.07 | 58.41 | 79.38 | 78.94 | 83.61 |
44 | 54.32 | 38.22 | 37.76 | 97.08 | 58.90 | 67.84 | 94.07 | 90.86 | 98.39 |
42 | 54.32 | 34.23 | 45.68 | 84.02 | 55.78 | 68.21 | 88.75 | 80.95 | 89.80 |
48 | 47.50 | 37.40 | 43.53 | 86.64 | 54.54 | 64.54 | 85.45 | 81.61 | 89.60 |
26 | 42.54 | 32.45 | 41.99 | 71.78 | 62.53 | 65.45 | 86.41 | 77.90 | 87.52 |
17 | 42.54 | 30.39 | 34.78 | 68.70 | 58.85 | 57.18 | 76.24 | 74.42 | 79.05 |
34 | 42.32 | 38.77 | 25.92 | 79.02 | 49.26 | 59.24 | 75.29 | 70.50 | 79.46 |
43 | 41.21 | 40.38 | 39.16 | 99.14 | 54.95 | 67.08 | 90.60 | 88.50 | 97.06 |
49 | 34.39 | 41.65 | 51.01 | 103.47 | 55.19 | 70.24 | 94.25 | 91.96 | 99.57 |
AS | 21.28 | 44.80 | 49.65 | 99.32 | 56.27 | 66.05 | 84.44 | 79.27 | 91.01 |
Model Index | #Params (M) | FGSM | PGD | ||||||||||||
Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | ||
IND | 106.38 | 66.78 | 7.88 | 41.82 | 3.91 | 3.76 | 73.20 | 79.15 | 153.94 | 16.57 | 94.23 | 5.47 | 4.40 | 178.51 | 179.32 |
190 | 105.03 | 53.91 | 11.82 | 34.50 | 5.91 | 6.65 | 56.94 | 58.89 | 105.63 | 37.71 | 116.13 | 6.74 | 6.36 | 141.97 | 157.73 |
348 | 105.02 | 53.04 | 12.33 | 40.33 | 6.35 | 5.90 | 59.07 | 64.92 | 115.81 | 38.49 | 120.48 | 9.71 | 6.22 | 153.14 | 166.49 |
200 | 104.81 | 53.40 | 11.13 | 36.40 | 5.03 | 6.38 | 56.96 | 61.64 | 104.28 | 36.37 | 113.03 | 5.74 | 7.49 | 140.94 | 155.51 |
352 | 104.80 | 51.95 | 11.90 | 35.91 | 6.08 | 5.80 | 55.51 | 59.68 | 102.05 | 36.28 | 116.72 | 8.21 | 5.88 | 138.80 | 156.63 |
469 | 104.57 | 53.07 | 12.95 | 40.12 | 9.45 | 7.10 | 58.58 | 64.07 | 105.87 | 37.13 | 122.31 | 16.36 | 9.68 | 144.30 | 160.39 |
358 | 103.68 | 53.60 | 12.46 | 36.65 | 6.14 | 5.72 | 56.95 | 60.86 | 107.96 | 38.20 | 123.47 | 8.41 | 6.16 | 148.42 | 164.93 |
481 | 103.46 | 57.14 | 13.82 | 37.24 | 11.58 | 7.93 | 60.54 | 65.79 | 121.96 | 36.53 | 125.33 | 21.67 | 14.79 | 158.55 | 175.44 |
191 | 98.21 | 51.06 | 11.31 | 33.85 | 4.79 | 5.30 | 54.51 | 56.15 | 101.20 | 41.67 | 121.75 | 5.97 | 5.29 | 136.27 | 151.80 |
959 | 96.86 | 51.77 | 12.40 | 37.55 | 8.17 | 8.17 | 59.28 | 63.09 | 113.43 | 37.06 | 127.30 | 5.54 | 7.52 | 149.44 | 170.95 |
958 | 83.75 | 56.76 | 11.56 | 38.07 | 8.34 | 6.16 | 63.94 | 66.90 | 116.17 | 37.81 | 132.04 | 6.21 | 5.85 | 153.90 | 177.06 |
1020 | 83.53 | 57.77 | 11.82 | 37.90 | 8.13 | 6.36 | 64.77 | 68.43 | 126.77 | 36.15 | 115.46 | 6.72 | 7.59 | 158.30 | 175.65 |
1043 | 75.59 | 59.98 | 11.12 | 40.90 | 8.55 | 12.49 | 67.54 | 73.56 | 119.75 | 32.93 | 116.46 | 8.93 | 29.73 | 155.25 | 170.59 |
1037 | 62.48 | 55.33 | 12.16 | 36.44 | 8.69 | 8.28 | 61.76 | 66.87 | 123.41 | 32.09 | 118.24 | 7.56 | 33.09 | 155.66 | 172.77 |
AS | 21.28 | 59.65 | 26.28 | 36.46 | 8.89 | 8.81 | 59.55 | 59.14 | 139.16 | 46.94 | 69.82 | 7.34 | 10.84 | 138.78 | 124.39 |
Model Index | #Params (M) | APGD | Auto-SAGE | |||||||
Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | Baseline | GB-MTA | ||
IND | 106.38 | 150.34 | 16.32 | 93.49 | 5.40 | 4.33 | 175.00 | 175.98 | 178.51 | 179.32 |
190 | 105.03 | 98.82 | 33.70 | 109.26 | 6.39 | 6.12 | 132.72 | 146.86 | 134.60 | 147.75 |
348 | 105.02 | 107.45 | 34.00 | 112.95 | 9.31 | 5.93 | 142.54 | 154.78 | 144.55 | 155.85 |
200 | 104.81 | 97.78 | 32.08 | 105.60 | 5.39 | 7.20 | 131.47 | 145.08 | 133.21 | 145.95 |
352 | 104.80 | 95.66 | 32.25 | 109.25 | 7.75 | 5.62 | 129.44 | 145.90 | 131.41 | 147.07 |
469 | 104.57 | 98.82 | 33.00 | 113.63 | 15.70 | 9.14 | 134.73 | 149.15 | 136.58 | 150.36 |
358 | 103.68 | 100.97 | 34.09 | 115.14 | 7.88 | 5.74 | 138.28 | 153.40 | 140.26 | 154.84 |
481 | 103.46 | 113.34 | 32.11 | 116.28 | 20.99 | 12.67 | 147.42 | 162.81 | 149.73 | 164.18 |
191 | 98.21 | 94.75 | 37.13 | 113.48 | 5.59 | 5.13 | 127.89 | 141.74 | 129.68 | 142.58 |
959 | 96.86 | 105.62 | 33.55 | 118.21 | 5.25 | 7.28 | 139.53 | 158.45 | 141.08 | 159.39 |
958 | 83.75 | 108.44 | 33.77 | 122.67 | 5.97 | 5.76 | 143.50 | 163.92 | 145.27 | 164.72 |
1020 | 83.53 | 117.91 | 31.86 | 107.94 | 6.30 | 7.10 | 147.29 | 162.43 | 149.50 | 163.90 |
1043 | 75.59 | 112.46 | 29.12 | 109.43 | 8.45 | 27.97 | 145.09 | 159.55 | 147.15 | 160.82 |
1037 | 62.48 | 115.08 | 28.25 | 110.43 | 7.28 | 29.00 | 145.04 | 160.81 | 151.71 | 165.31 |
AS | 21.28 | 136.60 | 46.05 | 69.10 | 7.37 | 10.73 | 135.97 | 122.22 | 138.78 | 124.39 |
Results of MTL Adversarial Transferability
Table 12 reports the numerical results for Figure 4 in the main paper, including the results for six multi-task models attacked by APGD Single-x variants. These models represent six levels of parameter sharing, ranging from all-share (AS/5L), where all five layers of the backbone model (5L) are shared, to independent models (IND/0L), where no layers are shared. We observe a distinct trend where the attack transferability decreases along with the reduction in levels of parameter sharing, irrespective of the specific attack method utilized.
Single-Segm | Single-Norm | Single-Dept | |
AS/5L | 0.46 | 0.53 | 0.15 |
4L | 0.26 | 0.45 | 0.15 |
3L | 0.22 | 0.33 | 0.12 |
2L | 0.16 | 0.28 | 0.09 |
1L | 0.12 | 0.31 | 0.08 |
IND/0L | 0.02 | 0.17 | 0.1 |
Results of Attack Performance on Adversarially Trained Multi-Task Models
As introduced in Section 4.4 in the main paper, we adopt the single task Friendly Adversarial Training (FAT) (Zhang et al. 2020) in the MTL context. Specifically, we modify the underlying adversarial samples generation process to the multi-task attacks we investigated in this paper, including naive multi-task adaptations Single and Total, and the proposed GB-MTA. The detailed adversarial training algorithm is described in Algorithm D, where for each mini-batch, we generate the adversarial samples with GB-MTA-PGD to train the model.
Algorithm 4 FAT GB-MTA
Table 13 reports the accuracy of multi-task models trained on NYUv2 with adversarial training. It includes metrics for all tasks and performance degradation (shown in columns containing ARP). We employ PGD-based naive multi-task attacks Single-x and Total as well as GB-MTA with to generate different adversarial data. It can be seen that after adversarial training, the average accuracy of the multi-task model is dropped by compared with training with clean data only (the “w/o AT” row).
Adv. Train | Semantic Seg. | Surface Normal Prediction | Depth Estimation | ARP | ||||||||||||
mIoU | Pixel Acc | ARP | Error | , within | ARP | Error | , within | ARP | ||||||||
Mean | Median | 11.25 | 22.5 | 30 | Abs. | Rel. | 1.25 | |||||||||
w/o AT | 25.88 | 58.05 | - | 17.28 | 15.11 | 36.45 | 71.32 | 84.91 | - | 0.55 | 0.21 | 64.61 | 89.95 | 97.39 | - | - |
Single-Segm | 14.53 | 46.89 | 31.54 | 17.59 | 16.18 | 29.37 | 73.14 | 87.22 | 4.60 | 0.68 | 0.25 | 53.84 | 83.77 | 95.00 | 13.46 | 16.53 |
Single-Norm | 17.22 | 48.22 | 25.20 | 17.96 | 16.07 | 27.33 | 74.48 | 87.42 | 5.58 | 0.66 | 0.26 | 56.18 | 84.59 | 95.16 | 12.30 | 14.36 |
Single-Dept | 18.10 | 49.33 | 22.54 | 17.72 | 16.06 | 30.74 | 72.32 | 86.66 | 4.20 | 0.66 | 0.28 | 55.82 | 83.80 | 94.67 | 14.51 | 13.75 |
Total | 14.99 | 47.50 | 30.13 | 17.72 | 15.99 | 30.26 | 72.98 | 86.55 | 4.21 | 0.66 | 0.28 | 56.23 | 84.16 | 94.74 | 13.86 | 16.07 |
GB-MTA | 15.67 | 47.67 | 28.67 | 17.54 | 15.93 | 29.93 | 74.11 | 87.18 | 3.64 | 0.67 | 0.28 | 55.46 | 83.40 | 94.39 | 15.64 | 15.98 |
The same phenomenon of decreased model accuracy and increased model robustness can also be observed in Tables 14 and 15, where FGSM-based multi-task attack variants are utilized when generating adversarial samples in adversarial training. Table 14 reports the accuracy of the adversarially trained multi-task models similar to Table 13, while Table 15 shows the ARP of multi-task models trained with and without FAT in Table 14 and attacked by multi-task attacks including FGSM variants of GB-MTA, Single, and Total with .
We first observe a accuracy drop with adversarial training from Table 14, that is, decreased model performance. Then we observe an increase in model robustness after adversarial training from the lower ARP after the attack reported in Table 15. For instance, when attacked by GB-MTA, ARP decreased from to . Furthermore, GB-MTA remains the most effective attack method, consistently outperforming the Single-x and Total baselines by up to 6.87%.
Adv. Train | Semantic Seg. | Surface Normal Prediction | Depth Estimation | ARP | ||||||||||||
mIoU | Pixel Acc | ARP | Error | , within | ARP | Error | , within | ARP | ||||||||
Mean | Median | 11.25 | 22.5 | 30 | Abs. | Rel. | 1.25 | |||||||||
w/o AT | 25.88 | 58.05 | - | 17.28 | 15.11 | 36.45 | 71.32 | 84.91 | - | 0.55 | 0.21 | 64.61 | 89.95 | 97.39 | - | - |
Single-Segm | 22.99 | 55.72 | 7.59 | 18.11 | 15.49 | 36.36 | 68.22 | 81.60 | 3.15 | 0.59 | 0.25 | 61.23 | 88.08 | 96.51 | 5.88 | 5.54 |
Single-Norm | 23.22 | 56.66 | 6.34 | 17.53 | 14.85 | 38.18 | 70.32 | 83.15 | 0.32 | 0.58 | 0.24 | 61.62 | 88.58 | 96.80 | 4.74 | 3.58 |
Single-Dept | 23.41 | 55.53 | 6.94 | 17.80 | 15.12 | 36.90 | 69.86 | 82.77 | 1.27 | 0.58 | 0.24 | 62.16 | 88.70 | 96.77 | 4.17 | 4.13 |
Total | 23.00 | 55.42 | 7.83 | 17.66 | 14.91 | 37.83 | 70.06 | 82.75 | 0.27 | 0.60 | 0.25 | 60.54 | 88.03 | 96.55 | 6.43 | 4.84 |
GB-MTA | 23.36 | 56.12 | 6.53 | 18.03 | 15.27 | 37.07 | 68.78 | 81.66 | 2.22 | 0.59 | 0.25 | 60.57 | 88.04 | 96.54 | 6.30 | 5.02 |
Adv. Train | Single-Segm | Single-Norm | Single-Dept | Total | GB-MTA |
w/o AT | 29.50 | 28.31 | 38.91 | 34.71 | 39.40 |
Single-Segm | 6.49 | 5.03 | 7.20 | 7.95 | 8.83 |
Single-Norm | 12.56 | 10.22 | 14.67 | 14.97 | 17.09 |
Single-Dept | 10.75 | 8.98 | 11.59 | 12.51 | 13.91 |
Total | 7.39 | 5.62 | 8.04 | 8.79 | 9.65 |
GB-MTA | 8.64 | 6.94 | 10.35 | 10.57 | 11.91 |
Visualization Results
Figure 6 visualizes adversarial samples generated given an image from (a) NYUv2 and (b) Taskonomy with different attack methods including Single-x attack, total attack, and the proposed GB-MTA and various attack strengths from 0 to 16. We show that along with the increase in the perturbation bound , the magnitude of the noise becomes larger and more visible. All attacks become more effective and thus lead to similar attack performance as shown in Figures 1 and 2 of the main paper.
Multi-Task Architectures with Model Index
We show the effectiveness of the proposed multi-task attack GB-MTA by conducting attack experiments on multi-task models with different sharing levels in the main paper. Tables 16 and 17 present the multi-task model architectures in the Layout format proposed by TreeMTL (Zhang, Liu, and Guan 2022b). A layout is a symbolized representation of a tree-structured multi-task architecture. For tasks and a backbone model with branching points, a layout , where is a list of task sets at the -th branching point. Task sets in are subsets of tasks and a task set means the set of tasks in sharing the -th block.
Model Index | #Params (M) | Layout |
IND | 63.83 | [[{0}, {2}, {1}], [{0}, {2}, {1}], [{0}, {2}, {1}], [{0}, {2}, {1}], [{0}, {2}, {1}]] |
9 | 63.6 | [[{1}, {0, 2}], [{1}, {0, 2}], [{1}, {2}, {0}], [{1}, {2}, {0}], [{1}, {2}, {0}]] |
14 | 63.6 | [[{2}, {0, 1}], [{2}, {0, 1}], [{2}, {1}, {0}], [{2}, {1}, {0}], [{2}, {1}, {0}]] |
4 | 63.6 | [[{1, 2}, {0}], [{1, 2}, {0}], [{0}, {2}, {1}], [{0}, {2}, {1}], [{0}, {2}, {1}]] |
23 | 63.59 | [[{0, 1, 2}], [{1}, {0, 2}], [{1}, {2}, {0}], [{1}, {2}, {0}], [{1}, {2}, {0}]] |
10 | 62.48 | [[{1}, {0, 2}], [{1}, {0, 2}], [{1}, {0, 2}], [{1}, {2}, {0}], [{1}, {2}, {0}]] |
5 | 62.48 | [[{1, 2}, {0}], [{1, 2}, {0}], [{1, 2}, {0}], [{0}, {2}, {1}], [{0}, {2}, {1}]] |
28 | 62.47 | [[{0, 1, 2}], [{2}, {0, 1}], [{2}, {0, 1}], [{2}, {1}, {0}], [{2}, {1}, {0}]] |
38 | 62.25 | [[{0, 1, 2}], [{0, 1, 2}], [{2}, {0, 1}], [{2}, {1}, {0}], [{2}, {1}, {0}]] |
41 | 61.13 | [[{0, 1, 2}], [{0, 1, 2}], [{0, 1, 2}], [{0}, {2}, {1}], [{0}, {2}, {1}]] |
33 | 55.43 | [[{0, 1, 2}], [{0, 1, 2}], [{1, 2}, {0}], [{1, 2}, {0}], [{0}, {2}, {1}]] |
39 | 55.43 | [[{0, 1, 2}], [{0, 1, 2}], [{2}, {0, 1}], [{2}, {0, 1}], [{2}, {1}, {0}]] |
42 | 54.32 | [[{0, 1, 2}], [{0, 1, 2}], [{0, 1, 2}], [{1, 2}, {0}], [{0}, {2}, {1}]] |
44 | 54.32 | [[{0, 1, 2}], [{0, 1, 2}], [{0, 1, 2}], [{1}, {0, 2}], [{1}, {2}, {0}]] |
48 | 47.5 | [[{0, 1, 2}], [{0, 1, 2}], [{0, 1, 2}], [{0, 1, 2}], [{0}, {2}, {1}]] |
26 | 42.54 | [[{0, 1, 2}], [{2}, {0, 1}], [{2}, {0, 1}], [{2}, {0, 1}], [{2}, {0, 1}]] |
17 | 42.54 | [[{0, 1, 2}], [{1, 2}, {0}], [{1, 2}, {0}], [{1, 2}, {0}], [{1, 2}, {0}]] |
34 | 42.32 | [[{0, 1, 2}], [{0, 1, 2}], [{1}, {0, 2}], [{1}, {0, 2}], [{1}, {0, 2}]] |
43 | 41.21 | [[{0, 1, 2}], [{0, 1, 2}], [{0, 1, 2}], [{1}, {0, 2}], [{1}, {0, 2}]] |
49 | 34.39 | [[{0, 1, 2}], [{0, 1, 2}], [{0, 1, 2}], [{0, 1, 2}], [{1}, {0, 2}]] |
AS | 21.28 | [[{0, 1, 2}], [{0, 1, 2}], [{0, 1, 2}], [{0, 1, 2}], [{0, 1, 2}]] |
Layout | ||||
IND | 106.38 | [[{0}, {1}, {2}, {4}, {3}], [{0}, {1}, {2}, {4}, {3}], [{0}, {1}, {2}, {4}, {3}], [{0}, {1}, {2}, {4}, {3}], [{0}, {1}, {2}, {4}, {3}]] | ||||
190 | 105.03 | [[{0}, {3}, {1, 2, 4}], [{0}, {3}, {4}, {1, 2}], [{0}, {3}, {4}, {1, 2}], [{0}, {3}, {4}, {2}, {1}], [{0}, {3}, {4}, {2}, {1}]] | ||||
348 | 105.02 | [[{1, 2, 3, 4}, {0}], [{0}, {1, 2}, {4}, {3}], [{0}, {1, 2}, {4}, {3}], [{0}, {4}, {3}, {2}, {1}], [{0}, {4}, {3}, {2}, {1}]] | ||||
200 | 104.81 | [[{0}, {3}, {1, 2, 4}], [{0}, {3}, {1, 2, 4}], [{0}, {3}, {4}, {1, 2}], [{0}, {3}, {4}, {2}, {1}], [{0}, {3}, {4}, {2}, {1}]] | ||||
469 | 104.57 | [[{1, 2, 3, 4}, {0}], [{1, 2, 3, 4}, {0}], [{0}, {1, 2}, {4}, {3}], [{0}, {4}, {3}, {2}, {1}], [{0}, {4}, {3}, {2}, {1}]] | ||||
358 | 103.68 | [[{1, 2, 3, 4}, {0}], [{0}, {3, 4}, {1, 2}], [{0}, {3, 4}, {1, 2}], [{0}, {4}, {3}, {2}, {1}], [{0}, {4}, {3}, {2}, {1}]] | ||||
481 | 103.46 | [[{1, 2, 3, 4}, {0}], [{1, 2, 3, 4}, {0}], [{0}, {3}, {1, 2, 4}], [{0}, {3}, {1}, {4}, {2}], [{0}, {3}, {1}, {4}, {2}]] | ||||
191 | 98.21 | [[{0}, {3}, {1, 2, 4}], [{0}, {3}, {4}, {1, 2}], [{0}, {3}, {4}, {1, 2}], [{0}, {3}, {4}, {1, 2}], [{0}, {3}, {4}, {2}, {1}]] | ||||
959 | 96.86 | [[{1, 2, 4}, {0, 3}], [{0, 3}, {4}, {1, 2}], [{0, 3}, {4}, {1, 2}], [{0, 3}, {4}, {2}, {1}], [{4}, {2}, {1}, {3}, {0}]] | ||||
958 | 83.75 | [[{1, 2, 4}, {0, 3}], [{0, 3}, {4}, {1, 2}], [{0, 3}, {4}, {1, 2}], [{0, 3}, {4}, {2}, {1}], [{0, 3}, {4}, {2}, {1}]] | ||||
1020 | 83.53 | [[{1, 2, 4}, {0, 3}], [{1, 2, 4}, {0, 3}], [{0, 3}, {4}, {1, 2}], [{0, 3}, {4}, {2}, {1}], [{0, 3}, {4}, {2}, {1}]] | ||||
1043 | 75.59 | [[{1, 2, 4}, {0, 3}], [{1, 2, 4}, {0, 3}], [{1, 2, 4}, {0, 3}], [{0, 3}, {2, 4}, {1}], [{2, 4}, {1}, {3}, {0}]] | ||||
1037 | 62.48 | [[{1, 2, 4}, {0, 3}], [{1, 2, 4}, {0, 3}], [{1, 2, 4}, {0, 3}], [{0, 3}, {2, 4}, {1}], [{0, 3}, {2, 4}, {1}]] | ||||
AS | 21.28 | [[{0, 1, 2, 3, 4}], [{0, 1, 2, 3, 4}], [{0, 1, 2, 3, 4}], [{0, 1, 2, 3, 4}], [{0, 1, 2, 3, 4}]] |